Orpheum Productions in Nevada is considering three mutually exclusive alternatives for lighting enhancements to one of its recording studios. Each enhancement will increase revenues by attracting directors who prefer this lighting style. The cash flow details, in thousands of dollars, for these enhancements are shown in the chart below. MARR is 10%/year.

End of Year Light Bar Sliding Spots Reflected Beam
0 -12000 -17000 -20000
1 2500 4000 0
2 2500 4000 4300
3 2500 4000 8600
4 2500 4000 12900
5 2500 4000 17200
6 2500 4000 21500

Based on an annual worth analysis, which alternative (if any) should be implemented?

Light Bar, Sliding Spots, Reflected Beam, None of the Above

Show the annual worth for each option (report your answers in thousands of dollars, as in the chart):

Light Bar: $


Sliding Spots: $


Reflected Beam: $
Answer #1

AW of Light bar = -12000 * (A/P, 10%,6) + 2500

= -12000 * 0.229607 + 2500

= -255.29

AW of sliding spots = -17000 * (A/P, 10%,6) + 4000

= -17000 * 0.229607 + 4000

= 96.67

AW of Reflected beam = -20000 * (A/P, 10%,6) + 2300 * (A/G, 10%,6)

= -20000 * 0.229607 + 2300 * 2.223557

= 522.03

As the AW of Reflected beam is highest it should be selected